OpenStudy (anonymous):

what line is perpendicular to y=-3/2x+2/3?

OpenStudy (anonymous):

y=-3/2 x+ 2/3 the slopes have to multiple to equal -1 so -3/2 *2/3 = -1 y= 2/3x +2/3

OpenStudy (anonymous):

wouldnt it be y=2/3x+3/2?

OpenStudy (anonymous):

it doesnt matter really it could be 2/3x+4 or 2/3x + 7/8 as long as the slopes multiply to equal -1 then the y intercept doesnt matter